# CoRaiS: Lightweight Real-Time Scheduler for Multi-Edge Cooperative Computing

### v1

@@ III-C System-level State Evaluation Model
III-C1 Service-oriented Performance Estimation
@@
By defining $\phi(x)$ and $\zeta$, we can focus on considering the primary performance factors of edges while formulating multi-edge cooperative scheduling problem and designing algorithms, and ignore the secondary factors (the heterogeneous hardware configuration and various resource allocation mechanism to services at edges). Furthermore, with $\phi_{i}(x)$, we can predict response times on any edge by inputting the request’s data size. This eliminates the need to analyze black box or white box code to obtain the necessary computation numbers for the request and restricts the computing device to the CPU.
III-C2 Service-oriented Workload Evaluation
@@
When evaluating the workload of edge $e_{i}$, we focus on the requests in $Q_{i}^{le}$ and $Q_{i}^{in}$, because only requests in these two queues will make use of the local resources of edge $e_{i}$. Three features are introduced to evaluate the workloads, including required computing time to complete requests in $Q_{i}^{le}$ (referred to as $c_{i}^{le}$), required data transmission time for requests in $Q_{i}^{in}$ (referred to as $t_{i}^{in}$), and required computing time to complete requests in $Q_{i}^{in}$ (referred to as $c_{i}^{in}$).
@@
$c_{i}^{le}$, $t_{i}^{in}$ and $c_{i}^{in}$ can be evaluated in customized mathematical approximation models. In this paper, we compute them by (1), (2) and (3) respectively. $r$ refers to a request. $\alpha_{r}$ refers to the data size of request $r$. $\beta_{r}^{i}$ denotes the distance between the source edge of $r$ and edge $e_{i}$. $C_{t}$ is a constant to represent the transmission speed for unit data through unit distance. In (1) and (3), we average the computation time required to complete all tasks across multiple copies. As for (2), we make two assumptions based on experience to predict required data transmission time. Firstly, the data transmission time is positively correlated with both data size and transmission distance. Secondly, edges can simultaneously receive data sent by other edges from different ports.
@@
Evaluating the workload before each scheduling operation allows us to obtain the real-time service capacity of edge $e_{i}$. This real-time system state knowledge is instrumental in making well-informed scheduling decisions.
$c_{i}^{le}=\frac{\sum_{r\in Q_{i}^{le}}\phi_{i}(\alpha_{r})}{\zeta_{i}}$ (1)
$t_{i}^{in}=\max\limits_{r\in Q_{i}^{in}}C_{t}\alpha_{r}\beta_{r}^{i}$ (2)
$c_{i}^{in}=\frac{\sum_{r\in Q_{i}^{in}}\phi_{i}(\alpha_{r})}{\zeta_{i}}$ (3)
III-D Problem Formulation
In this paper, the multi-edge cooperative scheduling problem in each $SR_{k}$ has similar formulation. We firstly define some parameters to formulate the $SR_{k}$ as $CoMEC=(\mathcal{E},\mathcal{W},\mathcal{V},\mathcal{P},\mathcal{I})$. =$\mathcal{E}=\{e_{q}\}_{q=1}^{Q}$ is a set of network edges where $s_{k}$ is deployed, $|\mathcal{E}|=Q$ is the number of edges. $\mathcal{W}$ is a $Q\times Q$ matrix that specifies the data transmission distance between any pair of edges, i.e. the entry of $\mathcal{W}$ at the $i^{th}$ row and $j^{th}$ column, denoted by $w_{ij}\in\mathbb{R_{+}}$, is the time cost to transmit one unit data from $e_{i}$ to $e_{j}$. For disconnected $e_{i}$ and $e_{j}$, $w_{ij}=+\infty$. $\mathcal{V}=\{\phi_{q}\}_{q=1}^{Q}$ is a set of functions and $\phi_{q}$ represents computation time estimation function of $e_{q}$ (explained in the section III-C1). $\mathcal{P}$ is a $Q$-dimensional vector as well. $p_{i}$ denotes the replica numbers of $s_{k}$ at $e_{i}$. $\mathcal{I}$ is a $N\times 3$ matrix that specifies the current workload evaluation of edges. The entry of $\mathcal{I}$ at $q^{th}$ row, denotes as $I_{q}$, represents the workload evaluation of $e_{q}$. $I_{q}=(c_{q}^{le},c_{q}^{in},t_{q}^{in})$ (the definitions are explained in the section III-C2).
The requests distributed in $SR_{k}$ can be modeled as $CoR=(\mathcal{R},\mathcal{L},\mathcal{F},\mathcal{D})$. =$\mathcal{R}=\{r_{z}\}_{z=1}^{Z}$ is a set of requests that require $s_{k}$, $|\mathcal{R}|=Z$ is the number of requests. $\mathcal{L}$ is a $Z\times Q$ matrix, the entry $l_{zq}$ represents whether $r_{z}$ is located at $e_{q}$ before scheduling, if yes, $l_{zq}=1$, else, $l_{zq}=0$ =$\mathcal{F}=\{f_{z}\}_{z=1}^{Z}$ is a set that records the size of input data for all $r_{z}\in\mathcal{R}$. $\mathcal{D}$ keeps the practical related data of requests.
Given $CoMEC$ and $CoR$, let $\mathcal{X}\in\{0,1\}^{Z\times Q}$ be a permutation matrix. For $\forall$ $x_{zq}\in\mathcal{X}$, $x_{zq}=1$ represents $r_{z}$ is dispatched to $e_{q}$. Then the objective function of multi-edge cooperative scheduling problem can be formulated as (4). The parameter $T_{q}$ in (4) denotes the required time to complete all requests that are scheduled to $e_{q}$ as $\mathcal{X}$. (4) indicates that the purpose is to get a $\mathcal{X}$ that can minimize the response time over all edges and all requests. $T_{q}$ in (4) can be computed by (5)-(9) with constraints (10) and (11). The definitions of primary notations in the formulation are summarized in Table I.
TABLE I: The definitions of primary notations
Notation
Definition
$r_{z}$
The request $z$
$e_{q}$
The edge $q$
$s_{k}$
The service $k$
$l_{zq}$
Whether $r_{z}$ is located at $e_{q}$. $l_{zq}\in\{0,1\}$
$x_{zq}$
Whether $r_{z}$ is dispatched to $e_{q}$. $x_{zq}\in\{0,1\}$
$f_{z}$
The input data size of $r_{z}$.
$\phi_{q}(f_{z})$
The computation time to deal with data of $r_{z}$ at $e_{q}$.
$p_{q}$
The replica number of $s_{k}$ at $e_{q}$
$c_{q}^{le}$
The computation time to complete backlogs in $Q_{q}^{le}$.
$c_{q}^{in}$
The computation time to complete backlogs in $Q_{q}^{in}$.
$w_{nq}$
The transmission distance $e_{n}$ and $e_{q}$.
$C_{t}$
@@ III-D Problem Formulation
A constant to represent the transmission speed for unit data through unit distance.
$t_{q}^{in}$
The remaining transmission time of backlogs in $Q_{q}^{in}$.
$[Z]$
{1, 2, …, Z}. Z denotes the number of requests.
$[Q]$
{1, 2, …, Q}. Q denotes the number of edges.
@@
$\mu_{q}=\frac{\sum_{z\in[Z]}l_{zq}x_{zq}\phi_{q}(f_{z})}{p_{q}}+c_{q}^{le},~{}~{}\forall q\in[Q]$ (5)
$\eta_{q}=\frac{\sum_{z\in[Z]}(1-l_{zq})x_{zq}\phi_{q}(f_{z})}{p_{q}}+c_{q}^{in},~{}~{}\forall q\in[Q]$ (6)
$\upsilon_{q}=\max\limits_{z\in[Z]}~{}x_{zq}f_{z}(\sum_{n\in[Q]}l_{zn}w_{nq}),~{}~{}\forall q\in[Q]$ (7)
$\kappa_{q}=\max(C_{t}\upsilon_{q},t_{q}^{in})$ (8)
$T_{q}=\max(\kappa_{q},\mu_{q})+\eta_{q},~{}~{}\forall q\in[Q]$ (9)
@@
To be specific, (5) predicts the required computing time to complete all requests that will be executed locally, including previous backlogs and new requests scheduled as $\mathcal{X}$. (6) evaluates the required computing time to complete all requests that are transferred from other edges to $e_{q}$, taking into account both backlogs and new requests scheduled as $\mathcal{X}$. (7) predicts the required longest transmission time for data of new requests that are transferred from other edges to $e_{q}$ as $\mathcal{X}$, since multiple edges transmitting data to the same edge in parallel is allowed in our multi-edge cooperative computing system. Considering that the data of backlogs may haven’t been transmitted to $e_{q}$, we design (8) to estimate the maximum transmission time for all requests that are transferred from other edges to $e_{q}$, including both backlogs and new requests scheduled as $\mathcal{X}$, through $\max$ operation over $C_{t}v_{q}$ and $t_{q}^{in}$. Furthermore, due to data transmission and local computing can run simultaneously, and the requests in $Q_{q}^{in}$ cannot be executed until all relevant data has been received by $e_{q}$, we design (9) to predict the required response time to complete all requests that are scheduled to $e_{q}$ as $\mathcal{X}$, including backlogs and new requests as well. Here $\max$ operation is used to select the case that consumes longer time. While computing a feasible $\mathcal{X}$, constrains (10) and (11) must be met to ensure all requests can be scheduled to only one available edge.
@@
IV Lightweight Real-time Scheduler Design
@@ IV-A Architecture of CoRaiS
Edge encoder: It is a module to embed edge features. At the beginning of training, input features of edges $\mathbf{f}$ are initialized based on the current edge states. To be specific, the input features $\mathbf{f}_{q}$ of $e_{q}$ include (i) coordinates $(x_{q},y_{q})$; (ii) the coefficients of $\phi_{q}$ and replica numbers $\zeta_{q}$; (iii) workload evaluation vector $I_{q}$, and $I_{q}=(c_{q}^{le},c_{q}^{in},t_{q}^{in})$. The encoder computes initial $d_{h}$-dimensional edge embeddings $\mathbf{f}_{q}^{(0)}$ through a learnable linear projection with parameters $\mathbf{W}_{e}$ and $\mathbf{b}_{e}$: $\mathbf{f}_{q}^{(0)}=\mathbf{W}_{e}\mathbf{f}_{q}+\mathbf{b}_{e}$. Then the embeddings are updated through $L$ attention layers, which is motivated by Transformer [34] and Attention Model [24]. Each layer consists of two sublayers: a multi-head attention layer (MHA) and an edge-wise fully connected layer (FC). A skip-connection and batch normalization (BN) are also used at each sublayer. The operations are formulated as (12), where $\mathbf{f}_{q}^{(l)}$ denotes the produced edge embedding by layer $l\in\{1,..,L\}$.
@@
Request Encoder: It has similar architecture and operations with edge encoder to embed request features. But the learnable parameters are different. The initial feature $h_{z}$ of $r_{z}$ includes (i) coordinates of the source edge of $r_{z}$; (ii) input data size of $r_{z}$. The initial features $\mathbf{h}_{z}^{(0)}$ are initialized to a $d_{r}$-dimensional embeddings by linear projection as (13). Following that, $K$ attention layers are used to update request embeddings. The operations are presented as (14), where $MHA_{r}^{k}$ and $FC_{r}^{k}$ are the parameters that are used to embed requests features. The operations of $\mathbf{MHA}_{r}^{k}$ are similar with them in $\mathbf{MHA}_{e}^{l}$.
@@
IV-B Training CoRaiS
The scheduling probability distribution $p_{\theta}(\pi)=\{a_{qz}\}_{q\in[Q],z\in[Z]}$ is produced by CoRaiS, from which we can sample a scheduling decision $\pi$. In order to train CoRaiS, we define the expectation of the maximum response time of all requests over edges as loss function: $\mathcal{L}(\theta|g)=\mathbb{E}_{p_{\theta}(\pi)}[L(\pi)]$. Given $\pi$, $L(\pi)=-\hat{u}^{\pi}$ is computed according to (18) and (19). Firstly, a local reward $u_{q}$ is estimated by (18), where $RL_{q}^{\pi}$ denotes the set of requests to be executed locally and $RT_{q}^{\pi}$ includes requests that are transferred from other edges to $e_{q}$, based on the decision $\pi$. $Y(\varpi_{m},e_{q})$ is a function to compute the transmission distance between the source edge $\varpi_{m}$ and the execution edge $e_{q}$ of $r_{m}$. The connotation of each equation in (18) is similar with (5)-(9). Then the global reward is formulated as (19).
$\begin{split}\mu_{q}=&\frac{\sum_{r_{m}\in RL_{q}^{\pi}}\phi_{q}(f_{m})}{\zeta_{q}}+c_{q}^{le}\\ \eta_{q}=&\frac{\sum_{r_{m}\in RT_{q}^{\pi}}\phi_{q}(f_{m})}{\zeta_{q}}+c_{q}^{in}\\ \kappa_{q}=&max\{\max\limits_{r_{m}\in RT_{q}^{\pi}}f_{m}~{}Y(\varpi_{m},e_{q}),~{}~{}t_{q}^{in}\}\\ u_{q}^{\pi}=&~{}-(max(\mu_{q},\kappa_{q})+\eta_{q})\end{split}$ (18)

### v2

@@ V-A System-level State Evaluation Model
V-A1 Service-oriented Performance Estimation
@@
By defining $\phi(x)$ and $\zeta$, we can focus on considering the primary performance factors of edges while formulating multi-edge cooperative scheduling problem and designing algorithms, and ignore the secondary factors (the heterogeneous hardware configuration and various resource allocation mechanism to services at edges). Furthermore, with $\phi_{i}(x)$, we can predict response times on any edges by inputting the request’s data size. This eliminates the need to analyze black box or white box code to obtain the necessary computation numbers for the request and restricts the computing device to the CPU.
V-A2 Service-oriented Workload Evaluation
@@
When evaluating the workload of edge $e_{i}$, we focus on the requests in $Q_{i}^{le}$ and $Q_{i}^{in}$, because only requests in these two queues will make use of the local resources of edge $e_{i}$. Five features are introduced to evaluate the workloads, including required computing time to complete requests in $Q_{i}^{le}$ (referred to as $c_{i}^{le}$), required data transmission time for requests in $Q_{i}^{in}$ (referred to as $t_{i}^{in}$), required computing time to complete requests in $Q_{i}^{in}$ (referred to as $c_{i}^{in}$), required results transmission time for requests in $Q_{i}^{le}$ (referred to as $b_{i}^{le}$) and required results transmission time for requests in $Q_{i}^{in}$ (referred to as $b_{i}^{in}$).
The five features can be evaluated in customized mathematical approximation models. In this paper, we compute them by (1)-(5) respectively. $r$ refers to a request. $f_{r}$ refers to the data size of request $r$ and $u_{r}$ denotes data volume of $r$’s results. $\varpi_{r}$, $e_{q}$ and $\delta_{r}$ represent the source edge, the execution edge and the predicted edge of request $r$, respectively. $\psi(e_{q},\varpi_{r})$ computes the distance between $e_{q}$ and $\varpi_{r}$, while $\psi(e_{q},\delta_{r})$ obtains the distance between $e_{q}$ and $\delta_{r}$. $C_{t}$ is a constant to represent the transmission speed for unit data through unit distance.
$c_{i}^{le}=\frac{\sum_{r\in Q_{i}^{le}}\phi_{i}(f_{r})}{\zeta_{i}}$ (1)
$t_{i}^{in}=\max\limits_{r\in Q_{i}^{in}}C_{t}f_{r}\psi(e_{q},\varpi_{r})$ (2)
$c_{i}^{in}=\frac{\sum_{r\in Q_{i}^{in}}\phi_{i}(f_{r})}{\zeta_{i}}$ (3)
$b_{i}^{le}=\max\limits_{r\in Q_{i}^{in}}C_{t}u_{r}\psi(e_{q},\delta_{r})$ (4)
$b_{i}^{in}=\max\limits_{r\in Q_{i}^{in}}C_{t}u_{r}\psi(e_{q},\delta_{r})$ (5)
@@
In (1) and (3), we average the computation time required to complete all tasks across multiple copies. As for (2), (4) and (5), we make two assumptions based on experience to predict required data transmission time. Firstly, the data transmission time is positively correlated with both data size and transmission distance. Secondly, edges can simultaneously receive data sent by other edges from different ports.
@@
Evaluating the workload before each scheduling operation allows us to obtain the real-time service capacity of edge $e_{q}$. This real-time system state knowledge is instrumental in making well-informed scheduling decisions.
TABLE II: The definitions of primary notations
Notation Definition
$SR_{k}$ The service $s_{k}$ oriented multi-edge cooperative subsystem.
$r_{z}$ The request $z$
$e_{q}$ The edge $q$
$s_{k}$ The service $k$
$l_{zq}$ Whether $r_{z}$ is located at $e_{q}$. $l_{zq}\in\{0,1\}$
$x_{zq}$ Whether $r_{z}$ is dispatched to $e_{q}$. $x_{zq}\in\{0,1\}$
$f_{z}$ The input data size of $r_{z}$.
$u_{z}$ The output data size of $r_{z}$.
$\phi_{q}(f_{z})$ The computation time to deal with data of $r_{z}$ at $e_{q}$.
$p_{q}$ The replica number of $s_{k}$ at $e_{q}$
$c_{q}^{le}$ The computation time to complete backlogs in $Q_{q}^{le}$.
$c_{q}^{in}$ The computation time to complete backlogs in $Q_{q}^{in}$.
$\psi(e_{q},e_{n})$ The transmission distance $e_{n}$ and $e_{q}$.
@@
$C_{t}$ A constant to represent the transmission speed for unit data through unit distance.
$t_{q}^{in}$ The remaining transmission time of backlogs in $Q_{q}^{in}$.
$[Z]$1, 2, …, Z}. Z denotes the number of requests.
$[Q]$1, 2, …, Q}. Q denotes the number of edges in one control region.
$[N]$1, 2, …, N}. N denotes the number of edges in the whole $SR_{k}$.
V-B Problem Formulation
Each service $s_{k}$ oriented subsystem $SR_{k}$ has multiple control regions. Each control region can be formulated as $CoMEC=(\mathcal{E},\mathcal{V},\mathcal{P},\mathcal{I})$. =$\mathcal{E}=\{e_{q}\}_{q=1}^{Q}$ is a set of network edges where $s_{k}$ is deployed, $|\mathcal{E}|=Q$ is the number of edges. $\mathcal{V}=\{\phi_{q}\}_{q=1}^{Q}$ is a set of functions and $\phi_{q}$ represents computation time estimation function of $e_{q}$ (explained in the section V-A1). $\mathcal{P}$ is a $Q$-dimensional vector as well. $p_{i}$ denotes the replica numbers of $s_{k}$ at $e_{i}$. $\mathcal{I}$ is a $N\times 3$ matrix that specifies the current workload evaluation of edges. The entry of $\mathcal{I}$ at $q^{th}$ row, denoted as $I_{q}$, represents the workload evaluation of $e_{q}$. $I_{q}=(c_{q}^{le},c_{q}^{in},t_{q}^{in})$ (the definitions are explained in the section V-A2). Additionally, we define $N$ to represent the number of edges in the whole $SR_{k}$. $\psi(e_{q},e_{n})$ is a function to obtain the data transmission distance between $e_{q}$ and $e_{n}$.
The requests distributed in an individual $CoMEC$ can be modeled as $CoR=(\mathcal{R},\mathcal{L},\mathcal{H},\mathcal{F},\mathcal{D})$. =$\mathcal{R}=\{r_{z}\}_{z=1}^{Z}$ is a set of requests that require $s_{k}$, $|\mathcal{R}|=Z$ is the number of requests. $\mathcal{L}$ is a $Z\times Q$ matrix, the entry $l_{zq}$ represents whether $r_{z}$ is located at $e_{q}$ before scheduling, if yes, $l_{zq}=1$, else, $l_{zq}=0$. $\mathcal{H}$ is a $Z\times N$ matrix, the entry $h_{zn}$ represents whether the client who submits $r_{z}$ will move to the client region of $e_{n}$ after $r_{z}$ is completed. If yes, $h_{zn}=1$, else, $h_{zn}=0$. =$\mathcal{F}=\{f_{z}\}_{z=1}^{Z}$ is a set that records the size of input data for all $r_{z}\in\mathcal{R}$. While $\mathcal{U}=\{u_{z}\}_{z=1}^{Z}$ is a set that records the size of results for all $r_{z}\in\mathcal{R}$. $\mathcal{D}$ keeps the practical related data of requests.
Given $CoMEC$ and $CoR$, let $\mathcal{X}\in\{0,1\}^{Z\times Q}$ be a permutation matrix. For $\forall$ $x_{zq}\in\mathcal{X}$, $x_{zq}=1$ represents $r_{z}$ is dispatched to $e_{q}$. Then the objective function of multi-edge cooperative scheduling problem can be formulated as (6). The parameter $T_{q}$ in (6) denotes the required time to complete all requests that are scheduled to $e_{q}$ as $\mathcal{X}$. (6) indicates that the purpose is to get a $\mathcal{X}$ that can minimize the response time over all edges and all requests. $T_{q}$ in (6) can be computed by (7)-(13) with constraints (14) and (15). The definitions of primary notations in the formulation are summarized in Table II.
@@ V-B Problem Formulation
$\mu_{q}=\frac{\sum_{z\in[Z]}l_{zq}x_{zq}\phi_{q}(f_{z})}{p_{q}}+c_{q}^{le}$ (7)
$\eta_{q}=\frac{\sum_{z\in[Z]}(1-l_{zq})x_{zq}\phi_{q}(f_{z})}{p_{q}}+c_{q}^{in}$ (8)
$\upsilon_{q}^{1}=\max\limits_{z\in[Z]}~{}x_{zq}f_{z}(\sum_{m\in[Q]}l_{zm}\psi(e_{m},e_{q}))$ (9)
$\kappa_{q}^{1}=\max(C_{t}\upsilon_{q}^{1},t_{q}^{in})$ (10)
$\upsilon_{q}^{2}=\max_{z\in[Z]}x_{zq}u_{z}(\sum_{n\in[N]}h_{zn}\psi(e_{q},e_{n}))$ (11)
$\kappa_{q}^{2}=\max(C_{t}\upsilon_{q}^{2},b_{q}^{in})$ (12)
$T_{q}=\max(\kappa_{q}^{1},\mu_{q})+\max(\eta_{q}+\kappa_{q}^{2},b_{q}^{in})$ (13)
@@
To be specific, (7) predicts the required computing time to complete all requests that will be executed locally, including previous backlogs and new requests scheduled as $\mathcal{X}$. (8) evaluates the required computing time to complete all requests that are transferred from other edges to $e_{q}$, taking into account both backlogs and new requests scheduled as $\mathcal{X}$. (9) predicts the required longest transmission time for data of new requests that are transferred from other edges to $e_{q}$ as $\mathcal{X}$, since multiple edges transmitting data to the same edge in parallel is allowed in our multi-edge cooperative computing system. Considering that the data of backlogs may haven’t been transmitted to $e_{q}$, we design (10) to estimate the maximum transmission time for all requests that are transferred from other edges to $e_{q}$, including both backlogs and new requests scheduled as $\mathcal{X}$, through $\max$ operation over $C_{t}v_{q}$ and $t_{q}^{in}$. (11) and (12) are used to estimate the longest transmission time from the execution edge $e_{q}$ to the predicted edge $e_{n}$ for requests that are processed at $e_{q}$, including new requests and backlogs. Furthermore, due to data transmission and local computing can run simultaneously, and the requests in $Q_{q}^{in}$ cannot be executed until all relevant data has been received by $e_{q}$, we design (13) to predict the required response time to complete all requests that are scheduled to $e_{q}$ as $\mathcal{X}$, including backlogs and new requests as well. Here $\max$ operation is used to select the case that consumes longer time. While computing a feasible $\mathcal{X}$, constrains (14) and (15) must be met to ensure all requests can be scheduled to only one available edge.
@@
VI Lightweight Real-time Scheduler Design
@@ VI-A Architecture of CoRaiS
Edge encoder: It is a module to embed edge features. At the beginning of training, input features of edges $\mathbf{f}$ are initialized based on the current edge states. To be specific, the input features $\mathbf{f}_{q}$ of $e_{q}$ include (i) coordinates $(x_{q},y_{q})$; (ii) the coefficients of $\phi_{q}$ and replica numbers $\zeta_{q}$; (iii) workload evaluation vector $I_{q}$, and $I_{q}=(c_{q}^{le},c_{q}^{in},t_{q}^{in},b_{q}^{le},b_{q}^{in})$. The encoder computes initial $d_{h}$-dimensional edge embeddings $\mathbf{f}_{q}^{(0)}$ through a learnable linear projection with parameters $\mathbf{W}_{e}$ and $\mathbf{b}_{e}$: $\mathbf{f}_{q}^{(0)}=\mathbf{W}_{e}\mathbf{f}_{q}+\mathbf{b}_{e}$. Then the embeddings are updated through $L$ attention layers, which is motivated by Transformer [40] and Attention Model [41]. Each layer consists of two sublayers: a multi-head attention layer (MHA) and an edge-wise fully connected layer (FC). A skip-connection and batch normalization (BN) are also used at each sublayer. The operations are formulated as (16), where $\mathbf{f}_{q}^{(l)}$ denotes the produced edge embedding by layer $l\in\{1,..,L\}$.
@@
Request Encoder: It has similar architecture and operations with edge encoder to embed request features. But the learnable parameters are different. The initial feature $h_{z}$ of $r_{z}$ includes (i) coordinates of the source edge of $r_{z}$; (ii) input data size of $r_{z}$; (iii) coordinates of the predicted edge of $r_{z}$; (iv) output data size of $r_{z}$. The initial features $\mathbf{h}_{z}^{(0)}$ are initialized to a $d_{r}$-dimensional embeddings by linear projection as (17). Following that, $K$ attention layers are used to update request embeddings. The operations are presented as (18), where $MHA_{r}^{k}$ and $FC_{r}^{k}$ are the parameters that are used to embed requests features. The operations of $\mathbf{MHA}_{r}^{k}$ are similar with them in $\mathbf{MHA}_{e}^{l}$.
@@
VI-B Training CoRaiS
The scheduling probability distribution $p_{\theta}(\pi)=\{a_{qz}\}_{q\in[Q],z\in[Z]}$ is produced by CoRaiS, from which we can sample a scheduling decision $\pi$. In order to train CoRaiS, we define the expectation of the maximum response time of all requests over edges as loss function: $\mathcal{L}(\theta|g)=\mathbb{E}_{p_{\theta}(\pi)}[L(\pi)]$. Given $\pi$, $L(\pi)=-\hat{u}^{\pi}$ is computed according to (22) and (23). Firstly, a local reward $u_{q}$ is estimated by (22), where $RL_{q}^{\pi}$ denotes the set of requests to be executed locally and $RT_{q}^{\pi}$ includes requests that are transferred from other edges to $e_{q}$, based on the decision $\pi$. $Y(\varpi_{m},e_{q})$ is a function to compute the transmission distance between the source edge $\varpi_{m}$ and the execution edge $e_{q}$ of $r_{m}$. $Y(\delta_{m},e_{q})$ obtains the transmission distance between the predicted edge $\delta_{m}$ and the execution edge $e_{q}$ of $r_{m}$. The connotation of each equation in (22) is similar with (7)-(13). Then the global reward is formulated as (23).
$\begin{split}\mu_{q}=&\frac{\sum_{r_{m}\in RL_{q}^{\pi}}\phi_{q}(f_{m})}{\zeta_{q}}+c_{q}^{le}\\ \eta_{q}=&\frac{\sum_{r_{m}\in RT_{q}^{\pi}}\phi_{q}(f_{m})}{\zeta_{q}}+c_{q}^{in}\\ \kappa_{q}=&\max\{\max\limits_{r_{m}\in RT_{q}^{\pi}}f_{m}~{}Y(\varpi_{m},e_{q}),~{}~{}t_{q}^{in}\}\\ \beta_{q}=&\max\{\max\limits_{r_{m}\in RT_{q}^{\pi}}u_{m}~{}Y(e_{q},\delta_{m}),b_{q}^{in}\}\\ u_{q}^{\pi}=&~{}-(\max(\mu_{q},\kappa_{q})+\max(\eta_{q}+\beta_{q},b_{q}^{le}))\end{split}$ (22)